OpenStudy (anonymous):

Solve the quadratic format. 2x^(2)-3x=1

OpenStudy (anonymous):

5/2 , 2

OpenStudy (anonymous):

is the exponent 2-3x or is it 2?

OpenStudy (anonymous):

its 2.

OpenStudy (anonymous):

2x^(2)-3x=1 2x^2 - 3x - 1 = 0 ->\[x = ( 3 \pm \sqrt{17} ) / 4\]